A new system of unit is devised in which unit length is 1 metre but the unit of mass and time are so chosen that c and G are of magnitude Unity ? What is the  new unit of mass in kg 
c - speed of light 
G - universal gravitational constant 
 
source :- " Problem for physics students " 
author :- KF Riley 

Grade:11

1 Answers

Aditya Gupta
4 years ago
G= 6.67408 × 10-11 m3 kg-1 s-2
and c= 299,792,458 m/s
let 1 sec= 299,792,458 sec’ where sec’ is new unit of time.
now since G has magnitude unity, 6.67408 × 10-11 = |sec^2|*|kg|
so |kg|= 6.67408 × 10-11 / |299,792,458|^2= p (say)
= so kg= p kg’ where kg’ is the new unit.
so kg’= 1/p kg= 134663530964*10^16
1.34663530964*10^27 kg
kindly approve :)